Three people put their money together to buy lottery tickets. The first person put in $20, the second person put in $25 and the third person put in $30. One of their tickets was a winning ticket. If they won $15 million, what was the first person's share of the jackpot?
Found 2 solutions by checkley77, Fombitz:
Answer by checkley77(12844)

20/(20+25+30)=20/75=.2667 or 26.67% of the amount bet.
.2667*15,000,000=$4,000,000 should be his share of the winnings.

Answer by Fombitz(32388)

They would share in the same ratio as they put in.
20:25:30
In lowest terms, that would be,
4:5:6
which happens to sum to 15.
.
.
.
The person who put in $20, would share 4/15ths of the jackpot, or $4,000,000.
The person who put in $25, would share 5/15ths of the jackpot, or $5,000,000.
The person who put in $30, would share 6/15ths of the jackpot, or $6,000,000.